OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3? A. y -1= - 3 (x - 4) B. Y+ 4 = 3(x+1) C. Y+ 4 = -3(x+1) D. Y - 4 = - 3(x - 1)

OpenStudy (anonymous):

? what lol

OpenStudy (midhun.madhu1987):

y - y1 = m(x - x1) x1 = 1 y1 = 4 m = -3

OpenStudy (midhun.madhu1987):

Plug in the values...

OpenStudy (anonymous):

so its d

OpenStudy (midhun.madhu1987):

yes.. Perfect :)